OpenStudy (anonymous):

A city’s population, P (in thousands), can be modeled by the equation P=130(1.03)^x, where x is the number of years after January 1, 2000. For what value of x does the model predict that the population of the city will be approximately 170,000? A)8 B)9 C)10 D)11

OpenStudy (anonymous):

Population is given in thousands. So, divide 170,000 by 1000 = 170 Now P=170 Substitute all the x values given in the options so that you get P=170. When x=9, P=170. Ans is C)9

OpenStudy (anonymous):

P = 170,000 170 000 = 130(1.03)^x log (170 000) = x log [130(1.03)] x = log (170 000) / log [130(1.03)] hmm... are you sure that the equation was: P=130(1.03)^x with x raised to a power?

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

thanks darling (: